# Human-Robot Cooperative Distribution Coupling for Hamiltonian-Constrained Social Navigation

### v1

@@ Abstract
Navigating in human-filled public spaces is a critical challenge for deploying autonomous robots in real-world environments. This paper introduces NaviDIFF, a novel Hamiltonian-constrained socially-aware navigation framework designed to address the complexities of human-robot interaction and socially-aware path planning. NaviDIFF integrates a port-Hamiltonian framework to model dynamic physical interactions and a diffusion model to manage uncertainty in human-robot cooperation. The framework leverages a spatial-temporal transformer to capture social and temporal dependencies, enabling more accurate pedestrian strategy predictions and port-Hamiltonian dynamics construction. Additionally, reinforcement learning from human feedback is employed to fine-tune robot policies, ensuring adaptation to human preferences and social norms. Extensive experiments demonstrate that NaviDIFF outperforms state-of-the-art methods in social navigation tasks, offering improved stability, efficiency, and adaptability [^1] The experimental videos and additional information about this work can be found at: https://sites.google.com/view/NaviDIFF..
@@ I Introduction
Recent research in social navigation has made notable progress by integrating techniques from deep reinforcement learning (DRL) [11, 12, 13], analytical mechanics [14, 15], and sociology [16, 17, 18]. However, current approaches, whether implicitly coupled [11, 12, 13, 19, 20] or decoupled [21, 22, 23, 24, 25, 26], still face two primary challenges: uncertainty in understanding human-robot cooperation and the lack of models that can fully capture the physical dynamics of HRI.
@@
Second, to manage the uncertainty of human-robot cooperation and capture temporal dependencies, we employ a diffusion model combined with a ST transformer. The diffusion neural network [46] stochastically infers the distribution of human-robot cooperation styles, addressing uncertainty in intent inference. Meanwhile, the ST transformer [11, 12, 47] captures both social and temporal dependencies in human behavior, enabling more accurate construction of PH mechanism and predictions of human trajectories.
@@
  • NaviDIFF is presented as an interpretable social navigation framework that ensures closed-loop stability and models environmental dynamics using the PH mechanism. It effectively manages diverse human-robot cooperative distributions through a diffusion neural network and captures social and temporal dependencies via a spatial-temporal (ST) transformer.
@@ II Related Works
Port-Hamilton Mechanism: The Hamiltonian approach, rooted in analytical mechanics, evolved from the principle of least action through the Euler-Lagrange equations and the Hamiltonian equations of motion. In contrast, the network approach, originating from electrical engineering, forms the backbone of systems theory. Port-Hamiltonian systems unify these perspectives by linking the network model’s interconnection structure (often referred to as the "generalized junction structure" in bond graph terminology) with a geometric framework defined by a Poisson or Dirac structure. This framework allows Hamiltonian dynamics to model total stored energy, energy dissipation, and system ports.
@@
Socially Aware Robot Navigation: Advancements in robotics and artificial intelligence [50] have led to the development of numerous socially-aware robot navigation planners [11, 12, 13, 20], evolving from early systems like MINERVA [51] and RHINO [52]. These planners can be categorized into decoupled and coupled strategies for HRI inference and path planning [53]. Decoupled methods [21, 22, 23, 24, 25, 26] often suffer from issues such as the "freezing robot" problem [16] and "reciprocal dance" [54] in dynamic environments due to the gap between map scanning and path planning.
Recently, coupled approaches leveraging game theory [55], geometry analysis [17], and variational methods [20] have been introduced to address these issues. RL-based planners have also achieved significant success by approximating human-like navigation through environmental dynamics representation. For instance, [27] addresses cooperative collision avoidance via pairwise agent interactions, while [56, 57] handle crowd-wise interactions. Other methods [19, 28] use ST graphs to explicitly model HRI, and transformer-based approaches [11] further improve performance.
@@ III Preliminary
Inspired by recent applications of the Hamiltonian mechanism in robotics [44, 45], the socially-aware robot navigation (SAN) task can also be modeled as a Hamiltonian-based multi-agent system [43], meeting the necessary physical criteria. In this context, we formulate the Hamiltonian-constrained socially-aware robot navigation task using the PH mechanism. The PH system offers an interpretable, energy-based model of physical environmental dynamics, ensuring closed-loop energy stability and capturing human-robot interaction patterns. The typical input-state-output port-Hamiltonian system can be defined as follows:
@@
where $\nabla_{x}:=\frac{\partial}{\partial x}$ denotes the gradient function, $\mathbf{H}(x)$ represents the system’s total energy, $\mathbf{J}(x)=-\mathbf{J}(x)^{\top}$ represents the system’s interactive energy, $\mathbf{R}(x)=\mathbf{R}(x)^{\top}$ denotes the system’s dissipative energy, and $\mathbf{G}(x)$ represents the external input. Here, $u$ is the robot policy input, $x$ is the robot’s state, and $y$ represents the conjugated variables of $u$. The PH system can be viewed as an input-state-output framework, where the input and output space are $m$-dimensional vector spaces, and the state space is an $n$-dimensional manifold. Additionally, a ST-graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is constructed using a spatial-temporal transformer network to model the interactions within the multi-agent system for the Hamiltonian mechanism. In this graph, each vertex represents an individual agent, and the spatial or temporal dependencies between them are abstracted as edges.
@@
The continuous procedure of the PH mechanism is adapted to the discretization of the DRL paradigm over a set of discrete time sequences ${\{t_{\alpha}\}}^{\infty}_{\alpha=0}$, where $(t_{\alpha+1}-t_{\alpha})=T>0$. The actions follow the zero-order hold input assumption $\{\forall t\in[t_{\alpha},t_{\alpha+1})~{};~{}a_{t}=u_{t_{\alpha}}\}$. We define the PH dependence $\dot{x}$ using the PH function $f_{PH}(\cdot)$, which is represented by a spatial-temporal transformer. The discretization of the continuous environmental dynamics is then abstracted into discrete states and actions, $s_{t}=x_{t_{\alpha}},a_{t}=u_{t_{\alpha}}$, as follows:
@@ IV-A Hamiltonian-Constrained Socially Aware Navigation
The SAN PH system, which satisfies Eq. 4, can be considered a passive system. Therefore, by employing the energy-balancing passivity-based control (EB-PBC) approach [59, 60], along with energy shaping and damping injection [61], we can derive the desired robot policy $u_{d}$. This achieves the closed-loop dynamics of the SAN system as follows:
@@
By aligning Eq. 3 and Eq. 5, the desired policy for the mobile robot can be defined as:
@@
Additionally, if the term $\mathbf{G}(x)$ is full-rank, the transformation of the SAN Hamiltonian mechanism from open-loop to close-loop dynamics is exact with respect to the policy input. However, for other agnostic systems, this transformation may not always be successfully addressed [62]. In this context, NaviDIFF not only optimizes the energy-balance PH policy using DRL, rather complex PDE or ODE inference, but also leverages a diffusion neural network to approximate the uncertainty in energy-based HRI dynamics.
@@ IV-B HRI State Representation and Cooperation Uncertainty
We extend the hybrid spatial-temporal transformer network from [11, 47] to infer latent spatial-temporal HRI dependencies for Hamiltonian term representation. The HRI spatial attentions $\mathbf{{Y}}_{\rm{S}}\in\mathbb{R}^{N\times T\times d}$ are generated by a spatial transformer to capture human-robot spatial correlations and interactions, while the agent temporal attentions $\mathbf{{Y}}_{\rm{T}}\in\mathbb{R}^{N\times T\times d}$ are captured by a temporal transformer, representing agents’ temporal predictability. Once the spatial and temporal HRI features are obtained, they are fused into a latent state $\mathbf{{Y}_{\rm{F}}}\in\mathbb{R}^{N\times T\times d}$ via multi-modal fusion transformer.
@@
$[\mathbf{R}_{\bf\theta}(\bf x)]_{ii}$ $=f_{fc}^{R}[\mathbf{F}_{\rm R}^{\rm ii}+{\sum}_{j\in[1,N]}(\mathbf{F}_{\rm R}^{\rm ij}+\mathbf{F}_{\rm R}^{\rm ji})];$
$[\mathbf{J}_{\bf\theta}(\bf x)]_{ij}$ $=f_{fc}^{J}[{\mathbf{F}_{\rm K}^{\rm ij}}-{\mathbf{F}_{\rm M}^{\rm ji}}],\quad\forall j\in[1,N];$
@@
Here, the system’s interactive energy matrix $\mathbf{J}_{\theta}(x)$ represents the physically-informed human-robot energy interaction or exchange, while system’s dissipative energy matrix $\mathbf{R}_{\theta}(x)$ captures the negative influence between agents The total energy of the social human-robot system is represented by the Hamiltonian term $\mathbf{H}_{\theta}(x)$. Once these three Hamiltonian components are generated based on the spatial-temporal feature representation of environmental dynamics from the spatial-temporal transformer, the PH-based robot policy is calculated using Eq. 7.
@@
Additionally, inspired by [13], a large-language model (LLM) [63] is used as human-robot interface to interpret human language commands. These commands are then converted into robot destinations and human preference configurations (e.g., preferred social distances and social norms) in $I_{C}$.
@@
After processing through the diffusion neural network, the final policy is sampled from $\mathcal{T}$ policies as $\hat{u}^{r}_{\theta}\sim\{\hat{u}^{0}_{\theta,\tau}~{}|~{}\tau\in[1,\mathcal{T}]\}$ and is parameterized to connect with the DRL actor. Notably, the gap between the continuous PH physical process and the discrete DRL decision-making process is addressed by Eq. 2.
@@ IV-C Training Procedure
The training procedure for NaviDIFF consists of the following key stages: 1) Proximal policy optimization (PPO)-based policy initialization [28, 50]; 2) RLHF reward learning [11, 64]; and 3) off-policy RL policy learning [11, 65]. Firstly, we employ a state-of-the-art on-policy RL algorithm, PPO [50], to pre-train an initial social robot policy $\pi_{0}$, optimizing the objective functions $\mathcal{L}_{\theta}$ and $\mathcal{L}_{\phi}$ in Eqs. 12 and 13. PPO updates policy gradients based on action advantages, $(A_{t}=A_{\pi}(a_{t},s_{t})=Q_{\phi}(a_{t},s_{t})-V_{\pi}(s_{t}))$, and minimizes a clipped-ratio loss over mini batches from the recent dataset (generated by $\pi$), with clip rate $\epsilon$, as follows:
@@ V-A Simulation Experiment Setup
Environment Configuration: We conduct the simulation experiments using the OpenAI Gym-based social robot navigation environment [11], with Hamiltonian-based environmental dynamics as defined in Eq. 2. Pedestrians are controlled by ORCA [67] or Social Force (SF) [68] personal policies, with the assumption that the mobile robot is invisible to pedestrians to ensure maximum social security training.

### v2

@@ Abstract
Navigating in human-filled public spaces is a critical challenge for deploying autonomous robots in real-world environments. This paper introduces NaviDIFF, a novel Hamiltonian-constrained socially-aware navigation framework designed to address the complexities of human-robot interaction and socially-aware path planning. NaviDIFF integrates a port-Hamiltonian framework to model dynamic physical interactions and a diffusion model to manage uncertainty in human-robot cooperation. The framework leverages a spatial-temporal transformer to capture social and temporal dependencies, enabling more accurate spatial-temporal environmental dynamics understanding and port-Hamiltonian physical interactive process construction. Additionally, reinforcement learning from human feedback is employed to fine-tune robot policies, ensuring adaptation to human preferences and social norms. Extensive experiments demonstrate that NaviDIFF outperforms state-of-the-art methods in social navigation tasks, offering improved stability, efficiency, and adaptability [^1] The experimental videos and additional information about this work can be found at: https://sites.google.com/view/NaviDIFF..
@@ I Introduction
Recent research in social navigation has made notable progress by integrating techniques from deep reinforcement learning (DRL) [11, 12], analytical mechanics [13, 14], and sociology [15, 16, 17, 18]. However, current approaches, whether implicitly coupled [11, 12, 19, 20] or decoupled [21, 22, 23, 24, 25, 26], still face two primary challenges: uncertainty in understanding human-robot cooperation and the lack of models that can fully capture the physical dynamics of HRI.
@@
Second, to manage the uncertainty of human-robot cooperation and capture temporal dependencies, we employ a diffusion model combined with a spatial-temporal (ST) transformer. The diffusion neural network [46] stochastically infers the distribution of human-robot cooperation styles, addressing uncertainty in intent inference. Meanwhile, the ST transformer [11, 12] captures both social and temporal dependencies in human behavior, enabling more accurate construction of PH mechanics and predictions of human trajectories.
@@
  • NaviDIFF is presented as an interpretable social navigation framework that ensures closed-loop stability and models environmental physical dynamics using the PH mechanics. It effectively manages diverse human-robot cooperative distributions through a diffusion neural network and captures social and temporal dependencies via a spatial-temporal (ST) transformer.
@@ II Related Works
Port-Hamilton Mechanics: The Hamiltonian approach, rooted in analytical mechanics, evolved from the principle of least action through the Euler-Lagrange equations and the Hamiltonian equations of motion. In contrast, the network approach, originating from electrical engineering, forms the backbone of systems theory. Port-Hamiltonian systems unify these perspectives by linking the network model’s interconnection structure (often referred to as the "generalized junction structure" in bond graph terminology) with a geometric framework defined by a Poisson or Dirac structure. This framework allows Hamiltonian dynamics to model total stored energy, energy dissipation, and system ports.
@@
Socially Aware Robot Navigation: Advancements in robotics and artificial intelligence [49] have led to the development of numerous socially-aware robot navigation planners [11, 12, 20], evolving from early systems like MINERVA [50] and RHINO [51]. These planners can be categorized into decoupled and coupled strategies for HRI inference and path planning [52]. Decoupled methods [21, 22, 23, 24, 25, 26] often suffer from issues such as the "freezing robot" problem [15] and "reciprocal dance" [53] in dynamic environments due to the gap between map scanning and path planning.
Recently, coupled approaches leveraging game theory [54], geometry analysis [16], and variational methods [20] have been introduced to address these issues. RL-based planners have also achieved significant success by approximating human-like navigation through environmental dynamics representation. For instance, [27] addresses cooperative collision avoidance via pairwise agent interactions, while [55, 56, 57] handle crowd-wise interactions. Other methods [19, 28] use ST graphs to explicitly model HRI, and transformer-based approaches [11] further improve performance.
@@ III Preliminary
Inspired by recent applications of the Hamiltonian mechanics in robotics [44, 45], the socially-aware robot navigation (SAN) task can also be modeled as a Hamiltonian-based multi-agent system [42], meeting the necessary physical criteria. In this context, we formulate the Hamiltonian-constrained socially-aware robot navigation task using the PH mechanics. The PH system offers an interpretable, energy-based model of physical environmental dynamics, ensuring closed-loop energy stability and capturing human-robot interaction patterns. The typical input-state-output port-Hamiltonian system can be defined as follows:
@@
where $\nabla_{x}:=\frac{\partial}{\partial x}$ denotes the gradient function, $\mathbf{H}(x)$ represents the system’s total energy, $\mathbf{J}(x)=-\mathbf{J}(x)^{\top}$ represents the system’s interactive energy, $\mathbf{R}(x)=\mathbf{R}(x)^{\top}$ denotes the system’s dissipative energy, and $\mathbf{G}(x)$ represents the external input. Here, $u$ is the robot policy input, $x$ is the robot’s state, and $y$ represents the conjugated variables of $u$. The PH system can be viewed as an input-state-output framework, where the input and output space are $m$-dimensional vector spaces, and the state space is an $n$-dimensional manifold. Additionally, a ST-graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is constructed using a spatial-temporal transformer network to model the interactions within the multi-agent system for the Hamiltonian mechanics. In this graph, each vertex represents an individual agent, and the spatial or temporal dependencies between them are abstracted as edges.
@@
The continuous procedure of the PH mechanics is adapted to the discretization of the DRL paradigm over a set of discrete time sequences ${\{t_{\alpha}\}}^{\infty}_{\alpha=0}$, where $(t_{\alpha+1}-t_{\alpha})=T>0$. The actions follow the zero-order hold input assumption $\{\forall t\in[t_{\alpha},t_{\alpha+1})~{};~{}a_{t}=u_{t_{\alpha}}\}$. We define the PH dependence $\dot{x}$ using the PH function $f_{PH}(\cdot)$, which is represented by a spatial-temporal transformer. The discretization of the continuous environmental dynamics is then abstracted into discrete states and actions, $s_{t}=x_{t_{\alpha}},a_{t}=u_{t_{\alpha}}$, as follows:
@@ IV-A Hamiltonian-Constrained Socially Aware Navigation
The SAN PH system, which satisfies Eq. (4), can be considered a passive system. Therefore, by employing the energy-balancing passivity-based control approach [59, 60], along with energy shaping and damping injection [61], we can derive the desired robot policy $u_{d}$. This achieves the closed-loop dynamics of the SAN system as follows:
@@
By aligning Eq. (3) and Eq. (5), the desired policy for the mobile robot can be defined as:
@@
Additionally, if the term $\mathbf{G}(x)$ is full-rank, the transformation of the SAN Hamiltonian mechanics from open-loop to close-loop dynamics is exact with respect to the policy input. However, for other agnostic systems, this transformation may not always be successfully addressed [62]. In this context, NaviDIFF not only optimizes the energy-balance PH policy using DRL, rather complex PDE or ODE inference, but also leverages a diffusion neural network to approximate the uncertainty in energy-based HRI dynamics.
@@ IV-B HRI State Representation and Cooperation Uncertainty
We extend the hybrid spatial-temporal transformer network from [11] to infer latent spatial-temporal HRI dependencies for Hamiltonian term representation. The HRI spatial attentions $\mathbf{{Y}}_{\rm{S}}\in\mathbb{R}^{N\times T\times d}$ are generated by a spatial transformer to capture human-robot spatial correlations and interactions, while the agent temporal attentions $\mathbf{{Y}}_{\rm{T}}\in\mathbb{R}^{N\times T\times d}$ are captured by a temporal transformer, representing agents’ temporal predictability. Once the spatial and temporal HRI features are obtained, they are fused into a latent state $\mathbf{{Y}_{\rm{F}}}\in\mathbb{R}^{N\times T\times d}$ via multi-modal fusion transformer.
@@
$[\mathbf{R}_{\theta}(\bf x)]_{ii}$ $=f_{fc}^{R}[\mathbf{F}_{\rm R}^{\rm ii}+{\sum}_{j\in[1,N]}(\mathbf{F}_{\rm R}^{\rm ij}+\mathbf{F}_{\rm R}^{\rm ji})];$
$[\mathbf{J}_{\theta}(\bf x)]_{ij}$ $=f_{fc}^{J}[{\mathbf{F}_{\rm K}^{\rm ij}}-{\mathbf{F}_{\rm M}^{\rm ji}}],\quad\forall j\in[1,N];$
@@
Here, the system’s interactive energy matrix $\mathbf{J}_{\theta}(x)$ represents the physically-informed human-robot energy interaction or exchange, while system’s dissipative energy matrix $\mathbf{R}_{\theta}(x)$ captures the negative influence between agents The total energy of the social human-robot system is represented by the Hamiltonian term $\mathbf{H}_{\theta}(x)$. Once these three Hamiltonian components are generated based on the spatial-temporal feature representation of environmental dynamics from the spatial-temporal transformer, the PH-based robot policy is calculated using Eq. (7).
@@
Additionally, a large-language model (LLM) [63] is used as human-robot interface to interpret human language commands. These commands are then converted into robot destinations and human preference configurations (e.g., preferred social distances and social norms) in $I_{C}$.
@@
After processing through the diffusion neural network, the final policy is sampled from $\mathcal{T}$ policies as $\hat{u}^{r}_{\theta}\sim\{\hat{u}^{0}_{\theta,\tau}~{}|~{}\tau\in[1,\mathcal{T}]\}$ and is parameterized to connect with the DRL actor. Notably, the gap between the continuous PH physical process and the discrete DRL decision-making process is addressed by Eq. (2).
@@ IV-C Training Procedure
The training procedure for NaviDIFF consists of the following key stages: 1) Proximal policy optimization (PPO)-based policy initialization [28, 49]; 2) RLHF reward learning [11, 64]; and 3) off-policy RL policy learning [11, 65]. Firstly, we employ a state-of-the-art on-policy RL algorithm, PPO [49], to pre-train an initial social robot policy $\pi_{0}$, optimizing the objective functions $\mathcal{L}_{\theta}$ and $\mathcal{L}_{\phi}$ in Eqs. (12) and (13). PPO updates policy gradients based on action advantages, $(A_{t}=A_{\pi}(a_{t},s_{t})=Q_{\phi}(a_{t},s_{t})-V_{\pi}(s_{t}))$, and minimizes a clipped-ratio loss over mini batches from the recent dataset (generated by $\pi$), with clip rate $\epsilon$, as follows:
@@ V-A Simulation Experiment Setup
Environment Configuration: We conduct the simulation experiments using the OpenAI Gym-based social robot navigation environment [11], with Hamiltonian-based environmental dynamics as defined in Eq. (2). Pedestrians are controlled by ORCA [67] or Social Force (SF) [68] personal policies, with the assumption that the mobile robot is invisible to pedestrians to ensure maximum social security training.